# ITEACNet: Inverted Teacher-studEnt seArch Conversation Network

### v2

@@ Introduction
Joint Optimization
To simultaneously train the model for both the target task and handling incomplete data, we jointly optimize the emotion loss, $L_{emo}$, and the imputation loss, $L_{rec}$.
@@
To enable the student model to better handle incomplete data, we simultaneously calculate the reconstruction loss for its last three layers of hidden features. We assign a progressively weighted adjustment to the model’s attention to each layer’s reconstruction loss:
@@
Experiments
Table 1: Ablation results for Emotion Context Change Encoding.
Dataset Local Global WAF
@@
Table 2: Ablation results for Inverted Teacher-Student framework. We report WAF scores(%), and higher WAF indicates better performance. The Decline metric is obtained by calculating the difference in performance between 0% and 70% missing rates.
@@
0.0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 Decline
IEMOCAP (four-class) GCNet TF - 73.89 72.86 71.52 70.53 69.15 68.65 67.62 65.79 8.10
@@
Table 3: Comparison of classification performance under Independent Model Evaluation (IME). We present WAF scores (%).
@@
0.0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 Average
IEMOCAP (four-class) CPM-Net 58.00 55.29 53.65 52.52 51.01 49.09 47.38 44.76 51.46
AE 74.82 71.36 67.40 62.02 57.24 50.56 43.04 39.86 58.29
CRA 76.26 71.28 67.34 62.24 57.04 49.86 43.22 38.56 58.23
MMIN 74.94 71.84 69.36 66.34 63.30 60.54 57.52 55.44 64.91
GCNet 78.36 77.48 77.34 76.22 75.14 73.80 71.88 71.38 75.20
ITEACNet 77.47 76.87 76.44 76.02 75.48 74.72 74.19 73.84 75.82
IEMOCAP (six-class) CPM-Net 41.05 37.33 36.22 35.73 35.11 33.64 32.26 31.25 35.32
AE 56.76 52.82 48.66 42.26 35.18 29.12 25.08 23.18 39.13
CRA 58.68 53.50 49.76 45.88 39.94 32.88 28.08 26.16 41.86
MMIN 56.96 53.94 51.46 48.42 45.60 42.82 40.18 37.84 47.15
GCNet 58.64 58.50 57.64 57.08 56.12 54.40 53.60 53.46 56.18
ITEACNet 59.02 58.69 58.53 57.95 57.55 56.75 56.03 55.94 57.56
CMUMOSI CPM-Net 71.90 68.91 71.12 70.59 67.95 65.88 64.02 61.79 67.77
AE 85.86 82.27 78.43 74.00 69.83 66.62 60.22 55.64 71.61
CRA 85.68 82.61 78.53 75.12 70.20 67.40 62.40 59.40 72.67
MMIN 85.20 81.91 78.22 74.60 70.14 67.72 64.04 61.53 72.92
GCNet 85.01 82.54 80.17 78.54 76.48 73.45 69.46 68.35 76.75
ITEACNet 84.90 83.35 81.04 77.90 76.99 71.94 70.27 68.93 76.92
CMUMOSEI CPM-Net 78.47 74.79 74.48 73.81 72.39 70.43 68.73 67.07 72.52
AE 86.66 84.37 82.58 80.57 78.80 76.43 74.26 72.81 79.56
CRA 86.48 84.19 82.25 80.12 78.55 75.85 74.07 72.46 79.25
MMIN 85.78 83.77 81.85 79.77 77.63 75.36 72.95 71.18 78.54
GCNet 87.12 86.50 85.50 84.53 83.55 82.44 80.27 80.20 83.76
ITEACNet 86.71 86.40 84.97 84.40 83.25 82.00 80.65 80.21 83.57
Evaluation Method
When evaluating the robustness of a model, it is crucial to consider its performance across various missing scenarios. Previous studies would define different scenarios, assuming that incidental factors hindering data collection would impact data at a fixed rate (Peng et al. 2020; Lian et al. 2023; Sun et al. 2023). Consequently, these works masked training and testing data at the same rate to simulate incomplete data scenarios, allowing them train the model and test its performance under this missing rate. This process will repeat with different missing rates to ensure experimental coverage of diverse missing scenarios. In this paper, we refer to this evaluation method as Independent Model Evaluation (IME).
Specifically, we mask the testing data at a fixed missing rate and then test the model’s performance under this missing rate without changing its weights. Subsequently, this process will repeat with different missing rates, and the average is taken as a comprehensive performance metric for the robustness.
Evaluation Metrics
Following (Lian et al. 2023), for CMU-MOSI and CMU-MOSEI, we use MAE as the loss function and select the model based on the best weighted average F1-score (WAF) on the validation set, reporting its WAF on the test set. For IEMOCAP, we adopt a 5-fold leave-one-session strategy and report the mean of the best WAF on the test sets. WAF is computed by excluding the samples with label 0.
Datasets
CMU-MOSI (Zadeh et al. 2016) contains 2,199 utterance-level video clips from 93 videos. Each video clip is annotated with an emotion intensity score ranging from -3 (strongly negative) to 3 (strongly positive).
CMU-MOSEI (Zadeh et al. 2018) is an iterative version of CMU-MOSI and includes 23,453 utterance-level video clips with the same annotation format as CMU-MOSI.
IEMOCAP (Busso et al. 2008) has two annotation formats: a four-class format and a six-class format. The four-class format consists of 5,531 utterance-level video clips, with each clip labeled as neutral, happy, sad, or angry. The six-class format includes 7,433 utterance-level video clips, with additional labels for excited and frustrated emotions.
Following (Lian et al. 2023)  [^1] https://github.com/zeroQiaoba/GCNet, we extract audio features using wav2vec (Schneider et al. 2019), text features using DeBERTa (Liu et al. 2019), and visual features using MA-Net (Zhao, Liu, and Wang 2021). Similar to (Lian et al. 2023), taking the case of 70% data missing as an example, when reading each mini-batch, each modality randomly resets 70% of the data to zero, but at least one modality is retained for each utterance.
Setup
Results
Emotion Context Change Encoding Performance
Recognizing the significance of context changes in Conversational Emotion Recognition (CER), we propose a novel encoding method called Emotion Context Change Encoding (ECCE). This method takes into account context change information from both local and global perspectives. To comprehensively analyze the performance of ECCE, we compare the performance of teacher models under different settings, as ECCE is not influenced by data missing. The results are presented in Table 1, where ”Local” denotes the encoding of local emotion context, and ”Global” signifies the encoding of global context changes. As the local emotion context information cannot be directly utilized, we are unable to conduct independent performance testing on it.
Figure 4 illustrates the complete ECCE, which takes into account these aspects effectively. It can be observed that there is a strong local correlation in the top-left and bottom-right corners, indicating that the model can perceive significant differences in the changing context between the two parts of the conversation data. Additionally, there are three more pronounced local correlation regions in the bottom-right, corresponding to three distinct local contextual patterns: frequent emotional changes, a local context of happiness, and a local context of excitement. These results effectively demonstrate the advantage of ECCE in considering changes in local context and context change.
Table 4: Comparison of classification performance under Unified Model Evaluation (UME). We present WAF scores (%).

### v3

@@ 3.4.3 Guidance of ECCE
3.5 Joint Optimization
To simultaneously train the model for both the target task and handling incomplete data, we jointly optimize the emotion loss, $L_{emo}$, and the distance loss, $L_{rec}$.
@@ 3.5 Joint Optimization
To enable the student model to better emulate the teacher model, we simultaneously calculate the vector distance loss for its last three layers of hidden features. We assign a progressively weighted adjustment to the model’s attention to each layer’s reconstruction loss:
@@
4 Experiments
4.1 Datasets
CMU-MOSI [48] contains 2,199 utterance-level video clips from 93 conversations. Each video clip is annotated with an emotion intensity score ranging from -3 (strongly negative) to 3 (strongly positive).
CMU-MOSEI [49] is an iterative version of CMU-MOSI and includes 23,453 utterance-level video clips from 3225 conversations with the same annotation format as CMU-MOSI.
IEMOCAP [50] has two annotation formats: a four-class format and a six-class format. The four-class format consists of 5,531 utterance-level video clips from 151 conversations, with each clip labeled as neutral, happy, sad, or angry. The six-class format includes 7,433 utterance-level video clips from 151 conversations, with additional labels for excited and frustrated emotions.
Following [13]  [^1] https://github.com/zeroQiaoba/GCNet, we extract audio features using wav2vec [51], text features using DeBERTa [52], and visual features using MA-Net [53]. Similar to [13], taking the case of 70% data missing as an example, when reading each mini-batch, each modality randomly resets 70% of the data to zero, but at least one modality is retained for each utterance.
4.2 Evaluation Metrics
Following [13], for CMU-MOSI and CMU-MOSEI, we use MAE as the loss function and select the model based on the best weighted average F1-score (WAF) on the validation set, reporting its WAF on the test set. For IEMOCAP, we adopt a 5-fold leave-one-session strategy and report the mean of the best WAF on the test sets. WAF is computed by excluding the samples with label 0.
4.3 Evaluation Method
When evaluating the robustness of a model, it is crucial to consider its performance across various missing scenarios. Previous studies would define different scenarios, assuming that incidental factors hindering data collection would impact data at a fixed rate [16, 13, 14]. Consequently, these works masked training and testing data at the same rate to simulate incomplete data scenarios, allowing them train the model and test its performance under this missing rate. This process will repeat with different missing rates to ensure experimental coverage of diverse missing scenarios. In this paper, we refer to this evaluation method as Independent Model Evaluation (IME).
Specifically, we mask the testing data at a fixed missing rate and then test the model’s performance under this missing rate without optimizing its weights. Subsequently, this process will repeat with different missing rates, and the average is taken as a comprehensive performance metric for the robustness.
4.4 Setup
4.5 Baselines
To evaluate the performance of our proposed ITEACH-Net, we consider the following state-of-the-art incomplete multimodal learning methods as baselines.
CPM-Net [36] projects all samples into a common space without considering missing patterns. It learns well-structured features by equipping with a clustering-like classification loss.
AE [54] employs an autoencoder to estimate missing data from partially observed inputs and jointly optimizes the autoencoder’s reconstruction loss with the classification loss of the downstream task.
CRA [41] combines a series of residual autoencoders into a cascading architecture for data imputation and jointly optimizes imputation and downstream tasks end-to-end.
MMIN [11] combines CRA with cycle consistency learning to predict the latent representations of missing modalities, exhibiting excellent performance under various missing conditions.
GCNet [13] employs graph neural networks to model speaker and context information separately, and uses LSTM to model temporal information, finally jointly optimizing for imputation and downstream tasks. This method outperforms others in incomplete learning for ERC.
5 Results
5.1 Emotion Context Changing Encoder Performance
Table 1: Ablation results for Emotion Context Change Encoder. The best performance is highlighted in bold.
[…]
@@ 5.1 Emotion Context Changing Encoder Performance
Recognizing the unique pattern of context changes in Emotion Recognition in Conversational (ERC), we propose a novel encoder called Emotion Context Change Encoder (ECCE). This method takes into account context change information from both local and global perspectives. To comprehensively analyze the performance of ECCE, we compare the performance of teacher models under different settings, as ECCE is not influenced by data missing. The results are presented in Table 1, where ”Local” denotes the encoding of local emotion context, and ”Global” signifies the encoding of global context changes. As the local emotion context information cannot be directly utilized, we are unable to conduct independent performance testing on it.
@@
Figure 5 illustrates the complete ECCE, which takes into account these aspects effectively. It can be observed that there is a strong local correlation in the top-left and bottom-right corners, indicating that the model can perceive significant differences in the changing context between the two parts of the conversation data. Additionally, there are three more pronounced local correlation regions in the bottom-right, corresponding to three distinct local contextual patterns: frequent emotional changes, a local context of happiness, and a local context of excitement. These results effectively demonstrate the advantage of ECCE in considering changes of emotion contexts in conversations.
Table 2: Ablation results for Inverted Teacher-Student framework. We report WAF scores(%), and higher WAF indicates better performance. The Decline metric is obtained by calculating the difference in performance between 0% and 70% missing rates. The best performance is highlighted in bold.
@@
[…]
IEMOCAP (four-class) PRE TF - 73.89 72.86 71.52 70.53 69.15 68.65 67.62 65.79 8.10
@@
5.2 Inverted Teacher-Student Performance
Dealing with incomplete data poses a significantly greater challenge than handling complete data. However, current methods do not equip models for handling incomplete data with more stronger abilities to follow the performance of models trained with complete data. Therefore, we propose the Inverted Teacher-Student traning framework, which decouples the two computations, and employs a student model with complex structures to handle incomplete data while emulating the performance of a teacher model with simpler structures trained with complete data. To comprehensively analyze the performance of ITS, we compare different training frameworks under different settings. The results are presented in Table 2, where the ”PRE” represents previous training framework, which uses only one model to handle complete and incomplete data during the training phase. The ”TS” training framework represents the vanilla Teacher-Student framework, where the teacher model is more complex than the student model. ”TF” is an abbreviation for Transformer. It’s worth noting that, in the TS training framework, the teacher model is equipped with four encoder layers, while the student model is equipped with three.
Firstly, the performance of the Transformer under PRE training framework is not satisfactory. It not only fails to achieve outstanding performance with complete data but also exhibits poor capability in handling incomplete data. The IST training framework brings about a noticeable improvement in performance. The student model, guided by a simpler teacher model, not only performs better with complete data, surpassing the baselines across all four tasks, but also demonstrates significant performance improvements in handling incomplete data. In transitioning from complete data to a severe missing scenario with a 70% data loss, the ITS training framework exhibits less performance degradation. Compared to the PRE training framework, ITS incurs a smaller loss of 2.88% in the four-class IEMOCAP task, 5.24% in the six-class IEMOCAP task, and 1.28% in the CMU-MOSI task. While it shows a slight disadvantage in the CMU-MOSEI task, the difference is negligible due to the similar performance under various missing rates for both frameworks. These results demonstrate the shortcomings of the PRE training framework, while the ITS training framework can better enhance the model’s robustness when dealing with incomplete data.
Furthermore, we propose the efficient augmentation of the student model’s complexity using Neural Architecture Search (NAS), aiming to enhance its capability in handling incomplete data. While the introduction of NAS doesn’t necessarily result in superior performance on complete data, it significantly enhances the model’s capability to handle incomplete data. The student model enhanced through NAS achieved the smallest performance decline across all four tasks, with reductions of 3.63% in the four-class IEMOCAP, 3.08% in the six-class IEMOCAP, 15.97% in CMU-MOSI, and 6.50% in CMU-MOSEI. This substantial surpasses the previous settings. These results effectively demonstrate that compared to models processing complete data, models handling incomplete data require stronger computational power to process the incomplete data more effectively. Even when trained alongside the same teacher, student models with greater computational capabilities will perform better.
(a) IEMOCAP(four-class)
(b) IEMOCAP(six-class)
(c) CMU-MOSI
(d) CMU-MOSEI
Figure 6: The vector distance between $H_{1}^{s}$ and $H_{1}$ for different student models under the same teacher configuration.
(a) IEMOCAP(four-class)
(b) IEMOCAP(six-class)
(c) CMU-MOSI
(d) CMU-MOSEI
@@ 5.2 Inverted Teacher-Student Performance
Figure 7: Comparison of classification performance under Independent Model Evaluation (IME). We present WAF scores (%).
5.3 Distance Loss Performance
Within the ITS framework, the student model primarily acquires knowledge from the teacher model through distance loss. To further investigate the benefits of distance loss on model robustness, we compare the vector mean absolute distance between the deep encodings $H^{s}_{1}$ of student models of varying complexities and the encoding $H_{1}$ of the teacher model. The results are presented in Figure 6, where the red lines signifie the student model under the NAS setup, and the blue lines represent the student model in the Transformer configuration, with the x-axis denoting the data missing rate.
It is observable that, with the exception of the CMU-MOSEI task, the student models without NAS enhancement exhibit superior performance in predicting the hidden layer representations of the teacher model. However, while the NAS-enhanced student models show competitive reconstruction performance on the CMU-MOSEI task at data missing rates of 10% to 40%, they significantly underperform under conditions of extreme data incompleteness.
Integrating the findings with the results presented in Table 2, it is evident that a more robust student model performs poorly in reconstruction tasks. Therefore, we can conclude that while the knowledge imparted by the teacher model within the ITS framework is crucial, it only serves as a guide for the student model. In this guidance process, the student model can gradually learn the correct way to handle incomplete data suitable for itself, rather than merely focusing on imitating the teacher model. This also reveals the difference between the ITS method and imputation-based methods, where reconstruction performance does not represent the model’s robustness.
5.4 ITEACH-Net Performance
(a) IEMOCAP(four-class)
(b) IEMOCAP(six-class)
Figure 8: Comparison of classification performance at different missing rates for models trained with 70% missing data.
5.4.1 Independent Model Evaluation Performance
To demonstrate the outstanding performance of ITEACH-Net in ERC tasks, we maintained the same evaluation method as previous works in Figure 7, enabling a fair comparison with the baselines.
Firstly, the Average metric reflect the overall performance of the model under different missing scenarios. ITEACH-Net surpasses the current state-of-the-art (sota) work, GCNet, by 0.52%, 1.38%, and 0.17% in the four-class IEMOCAP, six-class IEMOCAP, and CMU-MOSI tasks, respectively, achieving the new sota performance. Although ITEACH-Net’s performance in the CMU-MOSEI task is 0.19% lower than GCNet, considering the carefully design of GCNet’s network structure, this performance remains highly competitive.
Furthermore, observing ITEACH-Net’s performance under different missing rates, we find that despite GCNet utilizing complex network structures to additionally process speaker information in conversation, there is still no significant difference between our method and GCNet across all four tasks. Even in severe missing scenarios, such as at a 70% missing rate, our method demonstrates more pronounced advantages in the four-class IEMOCAP and six-class IEMOCAP tasks, achieving performance improvements of 2.46% and 2.48%, respectively. These results collectively demonstrate the excellent robustness of ITEACH-Net.
Table 3: The model parameters are implemented using the default settings provided in the official code of GCNet.
Model Module Parameters
AE[54] - 1.29M
CRA[41] - 2.11M
MMIN[11] - 3.87M
GCNet[13] - 34.14M
ITEACH-Net Teacher 2.19M
Student 3.88M
5.4.2 Unified Model Evaluation Performance
To more accurately demonstrate the model’s robustness in dynamic missing rate scenarios, we further employ the UME method to evaluate their performance. As previous works did not provide experimental results for this evaluation method, all results presented here are reproduced by us. The parameters for each model during the reproduction process are listed in Table 3.
Table 4: Comparison of classification performance under Random training strategy. The best performance is highlighted in bold.
@@ 5.4.2 Unified Model Evaluation Performance
[…]
IEMOCAP (four-class) AE[54] 57.55 56.43 55.26 53.59 51.50 49.53 47.10 45.16 52.02
CRA[41] 36.89 35.76 34.77 33.45 32.28 30.72 29.22 28.07 32.65
MMIN[11] 74.71 72.88 70.81 68.15 65.49 62.22 58.53 55.85 66.08
GCNet[13] 77.09 77.00 76.47 76.19 75.30 74.71 72.24 69.94 74.87
ITEACH-Net 77.61 77.23 76.46 75.97 75.12 73.99 72.40 71.73 75.06
IEMOCAP (six-class) AE[54] 28.10 27.82 27.51 27.49 27.00 26.27 25.78 25.05 26.88
CRA[41] 24.15 23.65 23.04 22.16 21.39 20.65 19.38 18.53 21.62
MMIN[11] 57.00 54.85 52.79 50.13 47.33 44.62 41.24 39.16 48.39
GCNet[13] 56.05 56.41 56.18 56.03 55.87 54.42 52.93 51.87 54.97
ITEACH-Net 59.90 59.32 58.61 57.86 57.01 56.36 55.29 54.01 57.30
CMUMOSI AE[54] 83.84 79.80 77.07 70.78 66.53 63.67 56.56 51.91 68.77
CRA[41] 76.84 73.87 70.75 66.82 62.96 58.93 55.75 50.90 64.60
MMIN[11] 84.57 80.74 76.63 74.69 69.57 64.28 59.68 57.05 70.90
GCNet[13] 84.37 82.62 79.64 77.86 75.50 73.12 71.34 67.11 76.44
ITEACH-Net 86.23 84.22 81.65 79.27 76.43 75.18 70.65 68.96 77.82
CMUMOSEI AE[54] 84.48 82.82 81.20 79.13 77.33 75.49 73.13 71.24 78.10
CRA[41] 75.48 74.06 72.56 70.95 69.18 67.24 64.84 63.56 69.73
MMIN[11] 85.75 83.12 80.55 77.52 75.10 71.44 68.17 65.38 75.88
GCNet[13] 86.84 86.06 85.25 84.48 83.80 82.13 80.66 79.64 83.61
ITEACH-Net 87.40 86.50 85.11 84.59 82.81 82.30 80.66 79.79 83.65
To demonstrate the limitations of the IME, we present in Figure 8 the performance of different methods at various missing rates in the four-class and six-class IEMOCAP tasks when the model is trained only on data with a 70% missing rate. The results show that AE and CRA almost completely lose their predictive ability. Although GCNet can perform well on incomplete data, due to the limitations of the training framework, it cannot handle complete data effectively, leading to a decline in model performance as the missing rate decreases. The MMIN method based on distillation maintains the ability to process complete data, but its performance drops significantly when dealing with incomplete data. In contrast, ITEACH-Net learns to handle complete data effectively, not only improving performance as the missing rate decreases but also maintaining stable performance across different missing rates. These results aptly reflect the necessity of testing model robustness using UME.
However, the optimal training approach to fully leverage the potential of the model within the UME method remains an open question. In this paper, we compare the performance of models under two different training strategies:
Random Strategy: To equip the model with the capability to process data with varying degrees of incompleteness, a straightforward approach involves blending training data with different missing rates. Under this training strategy, each batch of training data sampled by the model is randomly assigned a missing rate ranging from 0% to 70%, ensuring that each batch experiences a distinct level of incompleteness.
As illustrated in Table 4, it is observable that ITEACH-Net achieves optimal results across the Average metrics of four tasks, particularly demonstrating a significant performance advantage in the six-class IEMOCAP and CMU-MOSI tasks. Similarly, ITEACH-Net consistently outperformed all baselines across various missing rates in these two tasks, while in the four-class IEMOCAP and CMU-MOSEI tasks, it alternately excels or is comparable to GCNet.
Table 5: Comparison of classification performance under Progressive training strategy. The best performance is highlighted in bold.